• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name yez

    Log entry time 05:51:53 on February 12, 2011

    Entry number 346489


    keyword=L-HRS: replay results for run#1187 and problems-list

    Problems List:
    1, S1 TDC: some channels are a little noisy. Maybe because of EDTM
    signals. Vince told me there is a flag D.edtmbb to cut off EDTM but it is
    not working right now. Anyway, it looks like it is not a big deal
    2, S2m ADC: Multi-peaks, or empty channels are gone if Ped. substraction
    is off, but it seems that there are not many high-photon electron
    signals, or says, no big signals when high energy electrons pass through
    3, Cer TDC: Noisy. We should take a run without EDTM to see whether they
    are because of the signals.
    4, VDC Wires: It does not look as what we see during cosmic ray runs. Is
    that because of the Quadro-poles are off?
    5, Pion Rejectors: As usual, there are some missing channles,
    PR1#18,#19,#20,#33, PR2#16,#17,#32,#33,#34 are empty or low statistic,and
    PR1#21 is always very noisy. Some of them are in the two ADC connectors.
    I will check the connectors. However, Pion Rejctor works well so far to
    seperate electrons and pions.

    Figure 1

    Figure 2

    Figure 3

    Figure 4

    Figure 5

    Figure 6

    Figure 7

    Figure 8

    Figure 9

    Figure 10

    Figure 11

    Figure 12

    Figure 13

    Figure 14

    Figure 15

    Figure 16

    Figure 17

    Figure 18

    Figure 19

    Figure 20